Assume a competitive firm faces a market price of $100, a cost curve of C 1.00q2 30q 1,600 and a marginal cost curve of MC 2.00q 30 The firms profit maximizing output level is 35.00 units, the profit per unit is S-10.71, and total profit is: S-374.85. However, if the firm wanted to maximize the profit per unit, how much would it produce? It would produce units. (round your answer to two decimal places) If the firm produced this output level, what would be the profit? Its profit would be $ (round your answer to the nearest penny)
Answer #1

Under competitive market, equilibrium is attained where P = MC

100 = 2q + 30

2q = 70

q = 35 units

Profit = TR - TC = PQ - TC = 100(35) - [1(35)2 +30(35) +1600] = 3500 - 1225 - 1050 - 1600 = - 375